OpenStudy (smurfy14):

limit as x approaches 0. (cot(x))/(csc(x))

OpenStudy (lgbasallote):

\[\huge \frac{\cot x}{\csc x} \implies \frac{\frac{\cos x}{\sin x}}{\frac{1}{\sin x}} \implies \frac{\frac{\cos x}{\cancel{\sin x}}}{\frac{1}{\cancel{\sin x}}} \implies \cos x\] does that help?

OpenStudy (lgbasallote):

@smurfy14 you there?